OpenStudy (anonymous):

FInd all x, 0 less than or equal to x less than or equal to 2pi, such that 3cos^2x + 4sin x = 4. Please explain! :)

OpenStudy (anonymous):

start with the fact that \(\cos^2(x)=1-\sin^2(x)\) to get a quadratic equation in sine

OpenStudy (anonymous):

\[3\cos^2(x) + 4\sin(x) = 4\] \[3(1-\sin^2(x))+4\sin(x)-4=0\] \[3-3\sin^2(x)+4\sin(x)-4=0\] \[-3\sin^2(x)+4\sin(x)-1=0\] or maybe \[3\sin^2(x)-4\sin(x)+1=0\] then solve the quadratic equation

OpenStudy (anonymous):

what do you mean by "or maybe"?

OpenStudy (anonymous):

reverse all the signs in the quadratic

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

if you want to do that you can

OpenStudy (anonymous):

k, thanks:)